OpenStudy (xxlyh):

What is the value of y in the equation 9 + y = 8? 72 17 1 –1

alones (alones):

Okay so smplify both sides of the equation. y+9=8 then subtract 9 from both sides. y+9−9=8−9 and that would be your answer y=−1

OpenStudy (xxlyh):

thank you! that was my first choice but just wanted to make sure :) do you mind helping me with one more?

OpenStudy (aihberkhan):

An easier way is to isolate \(y\). You would do this by subtracting 9 from both sides and you would get \(y = -1\) :)

alones (alones):

Surely u can ;) make a new post also

OpenStudy (xxlyh):

oh ok ill make a new post my bad :')

alones (alones):

it's alright xP